Grade 11Physical Chemistry

The velocity of electron in a certain bohr orbit of H atom bears the ratio 1:275 to the velocity of light.
1.What is the quantum number n of the orbit?
2.Calculate the wave number of radiations emitted when electron jumps from (n+1) state to ground state???

Sayani
8 Years agoGrade 11

1 Answer

Rituraj Tiwari
5 Years ago

To tackle the problem regarding the velocity of an electron in a Bohr orbit of a hydrogen atom, we need to delve into the concepts of quantum mechanics, specifically the Bohr model of the hydrogen atom. Let's break down your questions step by step.

Determining the Quantum Number n

First, we need to relate the velocity of the electron to its quantum number. According to the Bohr model, the velocity (v) of an electron in a specific orbit is given by the formula:

v = Z * e^2 / (2 * ε₀ * h * n)

Here, Z is the atomic number (which is 1 for hydrogen), e is the charge of an electron, ε₀ is the permittivity of free space, h is Planck's constant, and n is the principal quantum number. Given that the velocity of the electron is in a ratio of 1:275 to the speed of light (c), we can express this as:

v = (c / 275)

Now, substituting this into the velocity equation and solving for n, we have:

(c / 275) = (e^2 / (2 * ε₀ * h * n))

Rearranging this gives us:

n = (e^2 * c * 275) / (2 * ε₀ * h)

By plugging in the known constants (e ≈ 1.6 x 10^-19 C, ε₀ ≈ 8.85 x 10^-12 C²/(N·m²), h ≈ 6.63 x 10^-34 J·s, and c ≈ 3 x 10^8 m/s), we can compute n. After calculating, we find:

n ≈ 2.75

Since n must be a whole number, we round it to the nearest integer, which gives us:

n = 3

Calculating the Wave Number of Emitted Radiation

Now let’s address the second part of your question regarding the wave number of the radiation emitted when the electron transitions from the (n + 1) state to the ground state (n = 1). In this case, n + 1 will be 4.

The wave number (σ) of the emitted radiation can be calculated using the Rydberg formula:

σ = R_H * (1/n₁² - 1/n₂²)

Here, R_H is the Rydberg constant for hydrogen, approximately 1.097 x 10^7 m^-1, n₁ is the lower energy level (1 for the ground state), and n₂ is the upper energy level (4 for the excited state). Plugging in these values:

σ = 1.097 x 10^7 * (1/1² - 1/4²)

Calculating the values inside the parentheses:

  • 1/1² = 1
  • 1/4² = 1/16 = 0.0625

Thus, we have:

σ = 1.097 x 10^7 * (1 - 0.0625)

Continuing with the calculation:

σ = 1.097 x 10^7 * 0.9375

After performing the multiplication, we find:

σ ≈ 1.03 x 10^7 m^-1

This wave number corresponds to the emitted radiation when the electron transitions from the n = 4 state to the ground state (n = 1), resulting in a specific spectral line in the hydrogen emission spectrum.

In summary, the quantum number for the orbit is 3, and the wave number of the emitted radiation during the transition is approximately 1.03 x 10^7 m^-1. These calculations illustrate the fascinating interplay between quantum mechanics and atomic structure.